find the equation of a line that is perpendicular to the line y=1/9x+7 and contains the point (-6,0)
Found 2 solutions by lwsshak3, sachi:
Answer by lwsshak3(11628)

find the equation of a line that is perpendicular to the line y=1/9x+7 and contains the point (-6,0)
***
standard form of equation for a straight line: y=mx+b, m=slope. b=y-intercept
given line: y=1/9x+7
slope,m=1/9
slope of line perpendicular to given line=-9 (negative reciprocal)
equation: y=-9x+b
solve for b using coordinates of given point (-6,0)
0=-9*-6+b
b=-54
Equation of perpendicular line:
y=-9x-54

Answer by sachi(548)

the slope of the given liney=1/9x+7=mx+c=1/9
so the slope of the line that is perpendicular to the line =-9
and contains the point (-6,0)
the equation of the line is (y-0)/(x-(-6))=-9
or y/(x+6)=-9
or y=-9(x+6)=-9x-54
or y+9x+54=0